# Follow the Flow: On Information Flow Across Textual Tokens in Text-to-Image Models

### v1

@@ Abstract
Text-to-Image (T2I) models often suffer from issues such as semantic leakage, incorrect feature binding, and omissions of key concepts in the generated image. This work studies these phenomena by looking into the role of information flow between textual token representations. To this end, we generate images by applying the diffusion component on a subset of contextual token representations in a given prompt and observe several interesting phenomena. First, in many cases, a word or multiword expression is fully represented by one or two tokens, while other tokens are redundant. For example, in “San Francisco’s Golden Gate Bridge”, the token “gate” alone captures the full expression. We demonstrate the redundancy of these tokens by removing them after textual encoding and generating an image from the resulting representation. Surprisingly, we find that this process not only maintains image generation performance, but also reduces errors by 21% compared to standard generation. We then show that information can also flow between different expressions in a sentence, which often leads to semantic leakage. Based on this observation, we propose a simple, training-free method to mitigate semantic leakage: replacing the leaked item’s representation after the textual encoding with its uncontextualized representation. Remarkably, this simple approach reduces semantic leakage by 85%. Overall, our work provides a comprehensive analysis of information flow across textual tokens in T2I models, offering both novel insights and practical benefits. [^1] We release our code at https://github.com/tokeron/lens
@@ 1 Introduction
Text-to-image (T2I) models consist of two main components: a text encoder and a diffusion model (Ho et al., 2020; Song & Ermon, 2019). The former processes the user’s prompt, transforming it into a representation that guides the latter in generating the image. While T2I models are powerful and widely used, misalignment issues frequently arise. Generated images often fail to capture key concepts from the user’s prompt, leading to catastrophic neglect (Chefer et al., 2023a), semantic leakage of properties between entities (Rassin et al., 2022), or incorrectly bind different concepts (Huang et al., 2023; Rassin et al., 2022). Prior work has attempted to address these issues by modifying the cross-attention mechanism between the diffusion model and the encoded textual tokens (Rassin et al., 2023; Chefer et al., 2023a; Dahary et al., 2024), under the implicit assumption that the text encoder accurately captures the user’s intent and that each textual token primarily represents the item it is intended to convey.
In this work, we challenge these assumptions and investigate how information flows within tokens during the textual encoding process, and what information is encoded in each textual token representation. [^2] Throughout this work, we study token representations at the output of the text encoder. For brevity, we sometimes omit the word “representation”. To this end, we analyze lexical items within prompts, where each lexical item is a single word or multiword expression that functions as a unit of meaning. Our analysis focuses on two complementary aspects. First, in-item information flow—how different sub-word token representations of the same lexical item encode and distribute information. For example, do all the token representations of the item “pelican” (“pe”, “lic”, and “an”) encode this lexical item (Fig. 1, top left). Second, inter-item information flow—how information flows between different lexical items (Feng et al., 2024; Feng & Steinhardt, 2024). For example, given the prompt “A businesswoman on a runway”, does the representation of “runway” change in the context of the word “businesswoman” (Fig. 1, bottom middle).
Figure 1: Our main findings. Left: Information within a lexical item is unevenly distributed across its token representations, with a few carrying most of the meaning. Middle: Context may cause incorrect information flow between items, distorting their representations and leading to unintended images. In this example, the representation of “runway” is distorted in the context of “a businesswoman”. Right: We propose a simple solution—encode the leaked item (“runway” in this case) separately and patch it into the representation of the prompt after textual encoding.
For in-item information flow (4), we adapt a causal intervention method (Toker et al., 2025; see Section 3.2) to generate images from individual token representations. Specifically, we first encode the entire sentence, and then generate an image from each individual contextualized token representation. The resulting image reflects the information that the diffusion model can reconstruct when conditioned on that single token’s contextual representation. Our experiments with FLUX (black-forest labs, 2024) show that, in most cases, each lexical item is fully represented by only one or two tokens, while its remaining tokens might be considered redundant (4.2). Finally, we demonstrate that our findings can be used to improve image–text alignment in image generation. To this end, we mask all the non-representative tokens (tokens that do not represent the lexical item) just before the diffusion, and generate images directly from the remaining representations. We find that this operation alone reduces image generation error by 21% compared to standard image generation.
For inter-item information flow (5), we adapt the same causal intervention method—this time to assess the information encoded in one lexical item with respect to others in the prompt. For instance, in the sentence a “A pool and a table”, we examine the information flow between the items “pool” and “table”. We find that in 89% of cases, each lexical item’s information is preserved and not influenced by others. However, in some cases, information flows between tokens. Interestingly, this flow can occur between lexical items that should not influence each other, for example, “table” influencing the representation of “pool” in the sentence above, changing pool’s representation to a “pool table”. We show that this leakage can be mitigated by patching the leaked lexical item’s representation with an uncontextualized representation of this item. Notably, this simple intervention prevents leakage in 85% of cases.
In summary, we analyze information flow in T2I models and uncover two key findings: (1) lexical items are often represented by just one or two tokens, and removing non-representative tokens can improve generation; (2) while different lexical items usually do not influence each other’s representations, when they do, it can happen between items that should not influence each other—leading to semantic leakage. Finally, based on our observation, we propose a simple method to mitigate semantic leakage, leading to a 85% reduction in leakage errors.
2 Related work
2.1 Interpretability
Interpretability in multimodal models.
Previous work on interpretability in T2I models analyzed the text encoder. Chefer et al. (2023b) examined the concepts encoded in the latent space of CLIP (Radford et al., 2021) and how different concepts relate to each other. Toker et al. (2024) studied the progression of textual representations through the layers of the text encoder, using the diffusion model as a lens into the encoding process. Another line of work explored the connection between the text encoder and the image generator (Tang et al., 2023). In this work we investigate how information flows between tokens within a single layer. We focus on the final layer, as it directly conditions the diffusion process.
Information flow in LLMs.
Interpretability research has examined how token information propagates via attention in LLMs (Vig & Belinkov, 2019; Clark et al., 2019). However, raw attention weights can be misleading (Pruthi et al., 2020; Jain & Wallace, 2019). To address this, methods such as Attention Rollout (Abnar & Zuidema, 2020) were developed to more reliably measure information flow through attention. Other studies have employed Patchscopes (Ghandeharioun et al., 2024) and the logit lens (nostalgebraist, 2020) to interpret the encoding processes in LLMs. These works demonstrate that modern language models go far beyond simple subword processing. For instance, Kaplan et al. (2025) showed that models perform an implicit detokenization process, effectively fusing subword tokens into coherent word-level representations. Feucht et al. (2024) reported an “erasure” effect, in which the contributions of earlier tokens are overwritten by later ones—suggesting the emergence of an internal lexicon. Another prominent direction is probing (Adi et al., 2016; Liu et al., 2019). For example, Zhang & Bowman (2018); Brunner et al. (2019) trained probes to classify, from a token’s intermediate representation, the identity of tokens a fixed number of positions away. However, these methods are not fully reliable, as probes can exploit spurious correlations (Belinkov, 2022). Moreover, these works typically assess whether information is present or flows between tokens, but not whether the next step in the pipeline (e.g., the LM head or decoder) actually uses this information. In our analysis, we focus on intervention methods that test whether the next component in the image generation pipeline can retrieve and use the relevant information.
2.2 Challenges and solutions in multimodal models.
The recent rise of T2I models has introduced several challenges, prompting various approaches to address them. This work shows that a better understanding of information flow between tokens can help mitigate many of these challenges.
Semantic leakage.
One notable challenge in T2I models is semantic leakage (Rassin et al., 2022). For example, given the prompt “a bat flying in a baseball stadium”, the generated image might include a baseball bat instead of a bat (the animal). A similar effect has been observed in language models, where the context influences generation even when it conceptually should not (Gonen et al., 2024).
Attribute binding.
Another issue is attribute binding, which concerns the correct association of attributes with entities mentioned in the prompt. For example, in the prompt “a yellow flamingo and a pink sunflower”, models must correctly bind “yellow” to “flamingo” and “pink” to “sunflower”. Several approaches have been proposed to address these challenges, including performing an optimization in the latent space to produce an attention pattern that is better aligned with the syntactic structure of the sentence (Rassin et al., 2023), and merging tokens to concentrate information into a single representation (Hu et al., 2024).
Catastrophic neglect.
A further prominent challenge is catastrophic neglect, where entities mentioned in the prompt are completely omitted from the generated image. Many works have attempted to mitigate this, one of the first being Attend and Excite (Chefer et al., 2023a), which optimizes the image latents to increase attention on the neglected entity tokens.
@@ 3.1 Intervention on the text encoder
Our goal in this paper is to evaluate the effect of information flow between textual token representations, which later condition the diffusion process. We are interested in analyzing how information is distributed across tokens within a lexical item (e.g., do all the tokens in “San Francisco’s Golden Gate Bridge” represent that item, or is the information more diffusely spread?). To do so, we must be able to interpret the information encoded in individual tokens. We are also interested in measuring how different lexical items influence one another. To this end, we seek to isolate a subset of tokens and evaluate their joint representation. For these purposes, we adjust text intervention in diffusion models (Toker et al., 2025), allowing us to generate images from arbitrary subsets of tokens within the encoded user prompt.
Given a prompt with $N$ tokens, $t_{1},t_{2},\ldots,t_{N}$, our goal is to isolate and interpret the information encoded by a subset of these tokens. Let $S\subset\{1,\ldots,N\}$ be the index set of selected tokens, where $1\leq|S|\leq N$. We begin by encoding the full prompt using the text encoder $E$, yielding the final hidden states $e_{1},\ldots,e_{N}$. Separately, we encode a sequence consisting entirely of pad tokens to obtain pad embeddings $p_{1},\ldots,p_{N}$. We then construct a patched prompt by replacing all hidden states outside $S$ with the corresponding pad embeddings:
$\tilde{t}_{i}=\begin{cases}e_{i}&\text{if }i\in S,\\ p_{i}&\text{otherwise}\end{cases}\quad\text{for }i=1,\ldots,N.$
The patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$ is then used to guide the diffusion model. Generating an image from this patched representation allows us to visualize and isolate the individual contributions of the selected tokens as interpreted by the diffusion model. To evaluate the information in the generated images, we use a vision-language model (VLM): For the inter-item information flow experiments, we assess whether the generated image represents the full lexical item, while for the in-item information flow experiments, we measure whether information from other lexical items is present in the generated image. See Section 3.2 for an illustration of the method applied to interpret individual subtoken representations.
@@ 3.2 Experimental setup
We experiment with FLUX-Schnell (black-forest labs, 2024)—a recent state-of-the-art T2I model. We also report similar findings on FLUX-dev in A.5.1. FLUX models employ T5-XXL (Raffel et al., 2019) as their text encoder, enabling bidirectional information flow between tokens. In A.5.2 we also include another model, SDXL-Turbo (Sauer et al., 2023), which uses a different text encoder—CLIP (Radford et al., 2021). Unlike T5-xxl, CLIP is a causal language model. See Appendix A.5.2 for the analysis of SDXL-Turbo.
@@
We use a subset of prompts from DrawBench (Saharia et al., 2022) and Parti (Yu et al., 2022b). In total, we use 1,053 prompts. For each prompt, we generate five images using different random seeds. To conduct the analysis in our paper, we first find the lexical items in each prompt. We use GPT-4o (OpenAI et al., 2024) to extract the lexical items, resulting in a total of 4,864 unique lexical items. See A.1 for the exact prompt we used and further technical details. We then use spaCy (Honnibal et al., 2020) to determine the part of speech of each lexical item, and retain only nouns, proper nouns, and adjectives, as these are typically concrete and can be identified in their visual representation. We end up with 3,891 unique lexical items and use them to evaluate both inter-item information flow and in-item information flow.
@@
To evaluate the content of generated images, we also employ Qwen2-VL-72B-Instruct (Wang et al., 2024), a model with strong general vision capabilities. We restrict our evaluation to binary (yes/no) questions. See A.1 for the exact prompt used and additional details.
Figure 2: –ֿ
Evaluating in-item information flow: Our proposed framework interprets the information flow within a lexical item. We generate images from each token comprising the lexical item (left-hand side) and analyze the resulting images using a VLM (right-hand side). In this example, we interpret the different tokens composing the item pelican and find that only one of them, lic, represents the concept of a pelican, while the other two tokens (“pe”, and “an”) do not.
4 In-item information flow
@@ 4.1 How is information distributed across tokens?
In this section, we explore how information is distributed across the token representations within a lexical item. We aim to answer two main questions: Do all tokens encode the same information? And is the information evenly distributed across tokens or concentrated in specific token representations? These questions are central, as many T2I applications (Chefer et al., 2023a; Rassin et al., 2023; Dahary et al., 2024) treat all tokens of a lexical item equally. Discovering a non-uniform distribution of information across tokens could improve application effectiveness by focusing on the most informative tokens. We apply our intervention method (Section 3.1) to the hidden representation of each token in the prompt. We then generate an image conditioned on each token’s contextualized representation. We use Qwen2-VL to assess whether an image represents the lexical item it is part of. We repeat this analysis for each lexical item, for each prompt in our combined dataset (Section 3.2). The results show that in 89% of the cases, at least one of the tokens fully represents the lexical item. Out of the remaining 11%, about 85% cannot be generated even from the full lexical item, suggesting that the model is unfamiliar with the concept. See A.4 for a deeper analysis of these cases. The remaining (1.3% of all lexical items) can be generated from the full lexical item, indicating that in these cases, the lexical item’s representation is distributed across multiple tokens. Focusing on the cases where at least one token represents the lexical item, we examine the number of representative tokens (those that result in an image that represents the item) and non-representative tokens (those that do not). We compare the proportion of representative tokens across different lexical items of different lengths. Our results (7 in A.3) show that typically one or two tokens represent the concept, while the remaining tokens are non-representative tokens. As the lexical item becomes longer, the number of non-representative tokens tokens increases. Focusing on these non-representative tokens, we find that they account for 52% of the tokens within the lexical items represented by least two tokens. In the next section, we examine the effect of removing non-representative tokens.
@@
4.2 Are all prompt tokens necessary?
Our analysis reveals that 52ֿof lexical item tokens are non-representative. In this section, we examine whether these tokens have any effect on image generation. To answer this question we apply our intervention method for each prompt, generating an image after masking all the non-representative tokens. We then use Qwen-VL to measure whether the image aligns with the prompt and compare it to the image generated without this intervention.

### v2

@@ Abstract
Text-to-image (T2I) models generate images by encoding text prompts into token representations, which then guide the diffusion process. While prior work has largely focused on improving alignment by refining the diffusion process, we focus on the textual encoding stage. Specifically, we investigate how semantic information is distributed across token representations within and between lexical items (i.e., words or expressions conveying a single concept) in the prompt. We analyze information flow at two levels: (1) in-item representation—whether individual tokens represent their lexical item, and (2) cross-item interaction—whether information flows across the tokens of different lexical items. We use patching techniques to uncover surprising encoding patterns. We find information is usually concentrated in only one or two of the item’s tokens—For example, in the item “San Francisco’s Golden Gate Bridge”, the token “Gate” sufficiently captures the entire expression while the other tokens could effectively be discarded. Lexical items also tend to remain isolated; for instance, the token “dog” encodes no visual information about “green” in the prompt “a green dog”. However, in some cases, items do influence each other’s representation, often leading to misinterpretations—e.g., in the prompt “a pool by a table”, the token pool represents a pool table after contextualization. Our findings highlight the critical role of token-level encoding in image generation, suggesting that misalignment issues may originate already during the textual encoding.
Follow the Flow: On Information Flow Across Textual Tokens in Text-to-Image Models
[…]
Figure 1: Our main findings. Left: Information within a lexical item is unevenly distributed across its tokens’ contextualized representations. In this example, one token carries the meaning of the entire item (e.g., lic represents a pelican, while pe and an do not). Right: Context may cause incorrect information flow between items. In the example, the representation of runway is influenced by the context of “a businesswoman” (vs. “businessman”).
@@ 1 Introduction
Text-to-image (T2I) models typically consist of two main components: a text encoder and a diffusion model (Ho et al., 2020; Song and Ermon, 2019). The former processes the user’s prompt, transforming it into a representation that guides the latter in generating the image. Though widely used, T2I models often exhibit prompt-image misalignment, where generated images fail to capture key concepts from the user’s prompt (Chefer et al., 2023a; Rassin et al., 2022; Huang et al., 2023a). Prior work has attempted to address these issues by modifying the diffusion stage, and particularly the cross-attention mechanism (Rassin et al., 2023; Chefer et al., 2023a; Dahary et al., 2024), under the implicit assumption that each textual token reliably encodes the item it is intended to convey. This raises two fundamental questions: First, is the information for an intended concept distributed evenly across the concept tokens after textual encoding, or concentrated in just a few tokens? Second, does each token exclusively encode that concept?
In this work, we examine this assumption and study how visual information is distributed across tokens during the textual encoding stage. We focus our analysis on lexical items—words or phrases that convey a single concept. We trace how item information is distributed both within the tokens of a single item (in-item), and across tokens of different items (cross-item) (see Fig.˜1 for examples of the different cases). Using a causal intervention framework (Toker et al., 2025), we systematically assess what information is encoded in each contextual token representation at the encoder’s output, using prompts from widely used T2I benchmarks as our evaluation setting (Section˜2.2). [^1] Throughout this work, we study token representations at the output of the text encoder. For brevity, we sometimes omit the word “representation”.
For in-item representation (Section˜3), we find that a lexical item’s meaning is typically concentrated in one or two representative tokens. Surprisingly, ablating the non-representative tokens reduces generation error by 21% in FLUX (black-forest labs, 2024). In contrast, when no representative tokens exist for a particular item, it is often omitted entirely from the image—an instance of item negligence (Chang et al., 2024a; Chefer et al., 2023a).
For cross-item interactions (Section˜4), we assess whether information flows between different lexical items in a prompt. We observe such cross-item flow in 11% of cases. Interestingly, the flow of information does not always follow sentence syntax, which can result in misrepresentation of items within specific contexts—especially with polysemous words. For instance, in the prompt “a pool by a table”, “pool” can wrongly suggest “table” refers to a billiard table.
In summary, we study the textual encoding and identify two main insights. First, Lexical items have concentrated representation—the meaning of the item is often concentrated in a single representative token. Second, while items in a sentence do not typically encode information regarding other items in the sentence, in some cases, often when polysemous are involved, semantic leakage (Dahary et al., 2024; Gonen et al., 2024; Rassin et al., 2022) occurs during the textual encoding phase, causing miss-interpretations of items.
Together, these insights suggest practical avenues for improving token-level intervention techniques, enhancing evaluation benchmarks, and guiding future encoder-aware T2I generation approaches.
@@ 2.1 Intervention on the Text Encoder
Our goal in this paper is to evaluate the effect of information flow between textual token representations, that subsequently condition the diffusion process. In Section˜3, we are interested in analyzing how information is distributed across tokens within a lexical item . To do so, we seek to interpret the information encoded in individual tokens. In Section˜4, we are interested in measuring how different lexical items influence one another. To this end, we seek to isolate a subset of tokens and evaluate their joint representation in the context of another item. For these purposes, we adjust a method for text intervention in diffusion models proposed by Toker et al. (2025), allowing us to generate images from arbitrary subsets of contextual token representations by masking the rest of the tokens in the sequence.
Figure 2: Evaluating in-item information flow. Our proposed framework interprets the information flow within a lexical item. We generate images from each token comprising the lexical item (left) and analyze them with a VLM (right). In this example, only the token lic represents the concept “pelican”, whereas pe and an do not.
Given a prompt with $N$ tokens $t_{1},t_{2},\ldots,t_{N}$, our goal is to isolate and interpret the information encoded by a subset of these tokens. Let $S\subset\{1,\ldots,N\}$ be the index set of selected tokens, where $0<|S|<N$. We begin by encoding the full prompt using the text encoder $E$, yielding the final hidden states $h_{1},\ldots,h_{N}$. Separately, we encode a sequence consisting entirely of pad tokens to obtain pad embeddings $p_{1},\ldots,p_{N}$. We then construct a patched prompt by replacing all hidden states outside $S$ with the corresponding pad embeddings:
$\tilde{t}_{i}=\begin{cases}h_{i}&\text{if }i\in S,\\ p_{i}&\text{otherwise}\end{cases}\quad\text{for }i=1,\ldots,N.$
The patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$ is then used to guide the diffusion model. Generating an image from this patched representation allows us to visualize and isolate the individual contributions of the selected tokens as interpreted by the diffusion model. To evaluate the information in the generated images, we use a vision-language model (VLM): For the in-item representation experiments, we assess whether the generated image represents the full lexical item, while for the cross-item interactions experiments, we measure whether information from other lexical items is present in the generated image. See Fig.˜2 for an illustration of the method on the  in-item representation  experiment.
@@ 2.2 Experimental Setup
Our main results are reported for FLUX-schnell (black-forest labs, 2024), a recent state-of-the-art T2I model. We repeat all experiments with FLUX-dev, SDXL-Turbo (Sauer et al., 2023) and SANA (Xie et al., 2024a). FLUX models employ T5-XXL (Raffel et al., 2019) as their text encoder, enabling bidirectional information flow between tokens. Unlike Flux, SDXL used CLIP (Radford et al., 2021), and SANA uses Gemma (Team et al., 2024). These text encoders are unidirectional, meaning the tokens are only influence by the tokens that came before. These changes cause some differences in the information flow between tokens, elaborated in Section˜6 and in Section˜A.3.2
@@
We use a subset of 1,053 prompts from DrawBench (Saharia et al., 2022) and PartiPrompts (Yu et al., 2022b) datasets, filtered to include 4–20 words prompts and exclude cases with added complexity unrelated to our focus (e.g., misspellings, written text, rare words). For each prompt, we generate five images using different random seeds. To extract the lexical items, we prompt GPT-4o (OpenAI et al., 2024), resulting in 4,864 unique items across prompts. See Section˜A.1.1 for further technical details. We then use spaCy (Honnibal et al., 2020) to determine the part of speech of each lexical item, and retain only nouns, proper nouns, and adjectives, as these are typically concrete and can be identified in their visual representation. We end up with 3,891 unique lexical items and use them to evaluate both in-item representation and cross-item interactions.
@@
To evaluate the content of generated images, we employ Qwen2-VL-72B-Instruct (Wang et al., 2024), a model with strong general vision capabilities. We restrict our evaluation to binary (yes/no) questions and use the model to assess prompt-image alignment and the presence of specific items in the image (see Section˜A.1.3 for additional details). To validate Qwen2-VL’s reliability, we also conducted a human evaluation on 100 randomly sampled cases, evenly split between in-item representation (Section˜3) and cross-item interactions (Section˜4) settings. The results indicate substantial agreement, with a Cohen’s Kappa values of 0.868 for in-item representation and 0.764 for cross-item interactions. Next, we determined the majority agreement among the human annotators and compared it to the model predictions. For cross-item interactions, the accuracy (F1) is 0.927 (0.933), and for cross-item interactions, the accuracy is 0.810 (0.740). These results suggest that Qwen2-VL predictions align well with human judgment for our tasks.
Acc. Relative %
Tokens
Removed
# Prompts Before After Unaffected Degraded Improved Accuracy $\Delta$ ($p$)
1 144 81.25 83.33 98.29 0.85 14.83 +2.08
2 98 82.65 88.78 100.00 0.00 35.27 +6.13
3 45 93.33 93.33 97.62 2.38 33.28 +0.00
4 24 87.50 91.67 100.00 0.00 33.36 +4.17
5+ 28 78.57 85.71 100.00 0.00 33.32 +7.14
Overall 339 83.48 87.02 98.90 0.71 25.00 +3.54
Table 1: Effect of removing non-representative tokens on prompt-level accuracy. The table reports: (i) accuracy before and after token removal, (ii) the percentage of originally successful prompts that remained successful (Unaffected) or became failures (Degraded), and (iii) the percentage of originally failed prompts that were corrected (Improved). Overall, removing non-representative tokens rarely harms generation and often improves it.
3 In-Item Representation
In this section, we analyze information flow within lexical items at the token level. We first show semantic information is unevenly distributed across tokens, with typically one or two representing the item (3.1). We then show non-representative tokens are largely redundant and may even harm generation (3.2). Finally, we examine items neglected from the generated image, tracing failures to poor textual encoding or gaps in visual grounding (3.3).
@@ 3.1 How is Information Distributed Across Tokens?
We begin by exploring how information is distributed across the token representations within a lexical item. We focus on two key questions: Do all tokens encode the same semantic information? And is the information evenly distributed across tokens, or rather concentrated in specific tokens? These issues are crucial as many T2I applications (Chefer et al., 2023a; Rassin et al., 2023; Dahary et al., 2024) treat all tokens of a lexical item in a given prompt equally. Uncovering an asymmetrical distribution of information could improve the effectiveness of such applications by focusing on the most informative tokens.
Given a prompt and a lexical item, we feed the prompt to the text encoder, obtaining contextualized token representations. We then identify the tokens comprising the item and apply our intervention method (Section˜2.1) to generate an image conditioned on each token’s contextualized representation (i.e., without letting other tokens in the prompt influence the diffusion process). Finally, we use Qwen2-VL to assess whether the image represents the overall lexical item it is part of. We define a representative token as one whose isolated representation results in an image containing its corresponding lexical item. Tokens that do not meet this criterion are considered non-representative. We repeat this analysis for each lexical item and for each prompt in our dataset (Section˜2.2).
Our results (Fig.˜3, bottom row) show that in most cases (89%), there is (at least) one representative token. Interestingly, in the remaining instances where no representative tokens exist, we find the item is also absent from an image generated from the full item (allowing all of the item’s tokens to guide generation) in 88% of instances.
Figure 3: Distribution of representative tokens per lexical item length. “Rep.” denotes the number of representative tokens; the bottom row aggregates over all lexical items. In most cases, one or two tokens sufficient to represent the entire lexical item. As item length increases, the number of non-representative tokens grows accordingly.
We next focus on instances where at least one token represents the lexical item, and examine the number of representative- and non-representative tokens across lexical items of different lengths, averaged across items. Our results (Fig.˜3, top rows) show that typically one or two tokens represent the concept, while the remaining tokens are non-representative. Further, as the token length of the lexical item becomes longer, the number of non-representative tokens increases (Fig.˜3, rightmost column). Inspecting all lexical items in our data composed of two or more tokens, these non-representative tokens account for 52% of their tokens. Therefore, we next examine the effect of removing non-representative tokens altogether.
@@
3.2 Tokens: Are They All Necessary?
We now examine whether non-representative tokens—tokens that do not result in an image representing the item—have any effect on image generation. To answer this question, we apply our intervention method to each prompt, this time generating an image after masking all the non-representative tokens. We use Qwen-VL to measure whether this image aligns with the prompt, and compare it to an image generated from the full prompt without intervention.